OpenStudy (anonymous):

DIMENSIONAL ANALYSIS 0.035mg=____cg

OpenStudy (anonymous):

what the monkey is a cg?

OpenStudy (anonymous):

all i can help with is to convert \(0.35\) mg to grams, move the decimal over 3 places to get \[0.000035\] grams

OpenStudy (anonymous):

centigram. I always solve these types of problems by multiplying by fractions. In this case\[0.035 \text{ mg} \times \frac{1 \text{cg} }{\_ \text{mg} } = \_\text{cg}\]

OpenStudy (anonymous):

oh is a cg one hundredth of a gram?

OpenStudy (anonymous):

Do you know how many milligrams there are in a centigram. Recall that milli means 0.001 and centi means 0.01

OpenStudy (anonymous):

thats the part i dont know

OpenStudy (anonymous):

i know how to multiply the fractions, but not what the conversions are

OpenStudy (anonymous):

mill is thousand, c is hundred that is all

OpenStudy (anonymous):

Well, if milli means 0.001 (1/1000) and centi means 0.01 (1/100) how many milli does it take to make one centi?

OpenStudy (anonymous):

\[\frac{ 0.01 }{ 0.001 } = ?\]